# BatteryLife: A Comprehensive Dataset and Benchmark for Battery Life Prediction

### v1

@@ 3.3. CyclePatch
where $\textbf{W}_{1}^{l}\in\mathbb{R}^{D_{2}\times D_{1}}$, $\textbf{W}_{2}^{l}\in\mathbb{R}^{D_{1}\times D_{2}}$, $\textbf{b}_{1}^{l}\in\mathbb{R}^{D_{1}}$, $\textbf{b}_{2}^{l}\in\mathbb{R}^{D_{2}}$, and $LN$ denotes layer normalization (Ba et al., 2016). $\textbf{z}^{l-1}_{i}$ is the input to the $l^{th}$ layer, and $\textbf{z}^{0}_{i}$ is initialized with $\hat{\textbf{X}}_{i}$. After process of all layers, the output $\textbf{z}^{L}_{i}$ from each cycle are concatenated:
@@ 4. EXPERIMENTS
Chemical System 15 63 1 1
Operation Temperature 8 3 2 4
Protocol 631 1 12 2
@@ 4.1. Experimental Setup
For most batteries in BatteryLife, we use the measured end-of-life as the life label. However, a few batteries remain above 80% SOH at the end of cycling. While some prior works (Ma et al., 2022; Zhang et al., 2024) use the last recorded cycle number as the life label, this can introduce large label noise. To mitigate this issue, we exclude batteries with a final SOH greater than $\lambda+2.5\%$. For those with a SOH in ($\lambda$, $\lambda+2.5\%$], we apply linear extrapolation to estimate the life label at $\lambda$ SOH. In this work, $\lambda=80\%$ for all datasets except CALB, where $\lambda=90\%$ is used, as most of the collected CALB batteries do not reach 80% SOH. Additionally, for CALB batteries, we use $Q_{1}$ as $Q_{0}$ in Equation 2 to calculate SOH in accordance with CALB regulations.
@@ 4.2. Evaluation Metrics
We employ two metrics to evaluate model performance: mean absolute percentage error (MAPE) and $\alpha$-accuracy (Roman et al., 2021). MAPE is widely adopted as the main evaluation metric for BLP (Severson et al., 2019; Zhang et al., 2024; Ma et al., 2022; Fei et al., 2023; Tang et al., 2022), and computed as:
@@ 4.5. Model Performance on Seen and Unseen Aging Conditions (RQ3)
Zn-ion MAPE $0.515_{\pm 0.067}$ $3.153_{\pm 0.187}$ $0.562_{\pm 0.125}$ $0.209_{\pm 0.011}$
15%-ACC $0.297_{\pm 0.084}$ $0.046_{\pm 0.028}$ $0.301_{\pm 0.093}$ $0.499_{\pm 0.028}$
Na-ion MAPE $0.255_{\pm 0.036}$ $4.115_{\pm 0.138}$ $0.286_{\pm 0.034}$ $0.215_{\pm 0.014}$
15%-ACC $0.406_{\pm 0.084}$ $0.000_{\pm 0.000}$ $0.341_{\pm 0.072}$ $0.484_{\pm 0.037}$
CALB MAPE $0.140_{\pm 0.009}$ $1.536_{\pm 0.147}$ $0.146_{\pm 0.051}$ $0.234_{\pm 0.005}$
15%-ACC $0.704_{\pm 0.053}$ $0.219_{\pm 0.109}$ $0.767_{\pm 0.088}$ $0.442_{\pm 0.012}$
@@ 4.6. Cross-domain Transferability (RQ4)
In real-world situations, batteries of new types are often fewer than batteries of old types so that it is important to develop a model that has cross-domain transferability. In this work, we evaluate the cross-domain transferability of models by investigating if a model pretrained on Li-ion dataset can be transferred to other domains, and report the results in Table 6. Specifically, we study the transferability of CPMLP which is the best model in Li-ion domain. Three transferability evaluations are considered: (1) Frozen: Directly applying the pretrained CPMLP on target domains. (2) Fine-tuning (FT): Fine-tune all parameters of the pretrained CPMLP. (3) Domain adaptation (DA): we implement a domain adaptation technique by minimizing maximum mean discrepancy (Han et al., 2022) between the target-domain and source-domain embeddings produced by the last hidden layer. By expectation, with transfer learning techniques, an effective transferrable model should outperform the best baseline that is trained using only target-domain training data.
From Table 6, we observe that frozen pretrained models perform poorly on target domains due to significant domain shifts across domains. Domain adaptation significantly improves model performance on Zn-ion and Na-ion datasets. Fine-tuned CPMLP slightly outperforms the baseline regarding the 15%-Acc on CALB dataset. However, there is still a distance to satisfactory BLP even if the transfer learning is leveraged. We also note that both domain adaptation and fine-tuning encounter challenges in certain cases, such as transferring from Li-ion dataset to CALB dataset regarding MAPE. Given the results of the cross-domain evaluations, BatteryLife still has considerable room for developing more transferrable models for BLP.
@@ A.1. Dataset Preprocessing
For all public datasets, the present study filtered out the batteries whose SOH is not degraded to 82.5%, and conducted the data statistics on the remained batteries. Note that BatteryArchive is an active website, and the data statistics in this paper were conducted on the available data until September 2024. As for BatteryLife, we further excluded the batteries whose life labels are no larger than 100. Additionally, we found that some batteries in the UL_PUR dataset contain sudden significant SOH drops without recovery. We conjecture that these drops are caused by certain equipment faults and thus removed these batteries. The statistics of BatteryLife were conducted on the remained batteries.
Furthermore, some outliers might exist in the cycling data. For the 7 datasets standardized by BatteryML (Zhang et al., 2024), we detected abnormal cycles by running a median filtering method on discharge capacity degradation trajectory following BatteryML (Zhang et al., 2024) and removed the detected outliers. As for other datasets, we first removed the rest performance tests (RPT) and formation cycles. The RPT cycles are removed because they comply with charge/discharge protocols that are different from the cycling protocols, and the RPT cycles are also generally detected as outliers by BatteryML (Zhang et al., 2024). Likewise, formation cycles are removed. Moreover, some cycles of sudden rise or drop in discharge capacities are manually removed. After data cleaning, each battery was stored in a data format consistent with BatteryML (Zhang et al., 2024).
@@ A.3. Overview of The Integrated Datasets
  • HNEI (Devie et al., 2018): In this dataset, there are fourteen 18650-format lithium-ion batteries (LG Chem, ICR18650 C2) whose cathode is a mixture of $\rm LiCoO_{2}$ and $\rm LiNi_{0.4}Co_{0.4}Mn_{0.2}O_{2}$, and the anode is graphite. The electrolyte composition is not disclosed, and the nominal capacity is 2.8 Ah for all batteries. There are 2 different protocols with a working temperature of 25 degree Celsius.
@@
  • SNL (Preger et al., 2020): This dataset consists of sixty-one batteries in the format of 18650 cylindrical batteries. The positive electrode covers $\rm LiFePO_{4}$, $\rm LiNi_{0.81}Co_{0.14}Al_{0.05}O_{2}$, and $\rm LiNi_{0.84}Mn_{0.06}Co_{0.1}O_{2}$, while the negative electrode is graphite. The electrolyte formula is unknown. There are 3 nominal capacities consisting of 1.1 Ah, 3.2 Ah, and 3.0 Ah. There are 23 different protocols in this dataset with 3 different working temperatures (15°C, 25°C and 35°C).
@@
[…]
@@
  • Tongji (Zhu et al., 2022): This dataset contains 130 commercial Li-ion batteries in a format of 18650 cylindrical batteries, covering 3 chemical systems consisting of 3 different positive electrodes, 2 different negative electrodes, and an identical electrolyte. This dataset includes 2 nominal capacities (3.5 Ah and 2.5 Ah), 3 operating temperatures, and 6 charge/discharge protocols.
  • Stanford (Cui et al., 2024): This dataset covers 41 batteries in the format of pouch batteries whose positive electrode is $\rm LiNi_{0.5}Mn_{0.3}Co_{0.2}O_{2}$, negative electrode is artificial graphite. The electrolyte is a mixture of $\rm LiPF_{6}$, ethylene carbonate, ethyl methyl carbonate, dimethyl carbonate, and vinylene carbonate. The nominal capacity of these batteries is 0.24 Ah. All batteries underwent different formation protocols but were cycled with an identical charge/discharge protocol under a working temperature of 25 degree Celsius.
  • XJTU (Wang et al., 2024c): There are 55 batteries in this dataset in the format of 18650 cylindrical batteries. For all batteries, the positive electrode is $\rm LiNi_{0.5}Co_{0.2}Mn_{0.3}O_{2}$ and the negative electrode is graphite. The electrolyte is unknown. The nominal capacity is 2 Ah. This dataset includes 2 different operation conditions with a working temperature of 20 degree Celsius.
  • ISU_ILCC (Li et al., 2024): This dataset comprises 251 502030-size Li-polymer batteries. The positive electrode is lithium nickel manganese cobalt oxide, and the negative electrode is graphite. The electrolyte remains unknown. And the nominal capacity is 0.25 Ah. There are 202 different charge/discharge protocols in this dataset, with 30 degree Celsius as the working temperature.
  • Zn-ion: This dataset consists of 95 zinc-ion batteries in the format of coin batteries. The positive electrode is $\rm MnO_{2}$ while the negative electrode is zinc metal. The electrolyte is unknown. The discharge capacity of the $10^{th}$ cycle was taken as the nominal capacity for each battery. All batteries underwent the same charge/discharge protocol at 25 degree Celsius.
@@ B.1. Further Details on Hyperparameters
In this subsection, we introduce in detail how we conducted the hyperparameter search. First, we split the data into training set, validation set, and testing set with a ratio of 6:2:2. For the Li-ion dataset, we used random seed 2021 to partition the dataset. For Zn-ion, Na-ion, and CALB datasets, considering the much smaller dataset size, we randomly generated three different splits using 42, 2021 and 2024 as the random seeds, respectively. For each model, we tried more than 20 different sets of hyperparameters and selected the hyperparameters that led to the lowest MAPE on the validation sets. The search ranges of hyperparameters are described as follows: batch size was chosen from [16, 32, 64, 128], learning rate was selected from [$5\times 10^{-4}$, $5\times 10^{-3}$, $1\times 10^{-3}$, dropout rate was chosen from [0, 0.05, 0.1], and embedding dimension was selected from [32, 64, 128, 256]. Additionally, for CyclePatch models, both the number of intra-cycle encoder layers and the number of inter-cycle encoder layers were tuned from 0 to 12. The number of CNN layers is tuned from 1 to 8.
@@ B.2. Implementation Details of Baselines
In this subsection, we introduce how we implement the baselines in detail. First, the cycling data are resampled into $\textbf{X}\in\mathbb{R}^{3\times 30000}$ according to the process introduced in Section 4.1. After that, the three variables in each cycle are normalized without destroying the inherent characteristic of cycling data:
@@
where $\textbf{p}_{i}\in\mathbb{R}^{300}$, $\textbf{v}_{i}\in\mathbb{R}^{300}$, $\textbf{s}_{i}\in\mathbb{R}^{300}$ represent the capacity, voltage, and current records in the $i^{th}$ cycle. The $Q_{nominal}$ denotes the nominal capacity. Let $\overline{\textbf{X}}_{i}$ represent the normalized cycling data of the $i^{th}$ cycle. Then, Transformer encoder, LSTM, BiLSTM, GRU, BiGRU, DLinear, PatchTST, Autoformer, iTransformer, and MICN can directly model $\overline{\textbf{X}}_{i}$ and produce prediction as follows:
@@
As for the DLinear, PatchTST, Autoformer, iTransformer, and MICN, we use the implementation in https://github.com/thuml/Time-Series-Library. Specifically, we slightly revise the classification implementation to make models usable for BLP.

### v3

@@ 3.3. CyclePatch
where $\textbf{W}_{1}^{l}\in\mathbb{R}^{D_{2}\times D_{1}}$, $\textbf{W}_{2}^{l}\in\mathbb{R}^{D_{1}\times D_{2}}$, $\textbf{b}_{1}^{l}\in\mathbb{R}^{D_{2}}$, $\textbf{b}_{2}^{l}\in\mathbb{R}^{D_{1}}$, and $LN$ denotes layer normalization (Ba et al., 2016). $\textbf{z}^{l-1}_{i}$ is the input to the $l^{th}$ layer, and $\textbf{z}^{0}_{i}$ is initialized with $\hat{\textbf{X}}_{i}$. After process of all layers, the output $\textbf{z}^{L}_{i}$ from each cycle are concatenated:
@@ 4. EXPERIMENTS
Chemical System 13 45 1 1
Operation Temperature 8 3 1 4
Protocol 406 1 12 2
@@
[…]
[error curve]The performance of the top-three models with an increasing number of usable cycles.
@@ 4.1. Experimental Setup
For most batteries in BatteryLife, we use the measured end-of-life as the life label. However, a few batteries remain above 80% SOH at the end of cycling. While some prior works (Ma et al., 2022; Zhang et al., 2023a) use the last recorded cycle number as the life label, this can introduce large label noise. To mitigate this issue, we exclude batteries with a final SOH greater than $\lambda+2.5\%$. For those with a SOH in ($\lambda$, $\lambda+2.5\%$], we apply linear extrapolation to estimate the life label at $\lambda$ SOH. In this work, $\lambda=80\%$ for all datasets except CALB, where $\lambda=90\%$ is used, as most of the collected CALB batteries do not reach 80% SOH. Additionally, for CALB batteries, we use $Q_{1}$ as $Q_{0}$ in Equation 2 to calculate SOH in accordance with CALB regulations.
@@ 4.2. Evaluation Metrics
We employ two metrics to evaluate model performance: mean absolute percentage error (MAPE) and $\alpha$-accuracy (Roman et al., 2021). MAPE is widely adopted as the main evaluation metric for BLP (Severson et al., 2019; Zhang et al., 2023a; Ma et al., 2022; Fei et al., 2023; Tang et al., 2022), and computed as:
@@ 4.5. Model Performance on Seen and Unseen Aging Conditions (RQ3)
Zn-ion MAPE $0.515_{\pm 0.067}$ $3.153_{\pm 0.187}$ $0.562_{\pm 0.125}$ $0.752_{\pm 0.102}$
15%-ACC $0.297_{\pm 0.084}$ $0.046_{\pm 0.028}$ $0.301_{\pm 0.093}$ $0.070_{\pm 0.033}$
Na-ion MAPE $0.255_{\pm 0.036}$ $4.115_{\pm 0.138}$ $0.286_{\pm 0.034}$ $0.641_{\pm 0.121}$
15%-ACC $0.406_{\pm 0.084}$ $0.000_{\pm 0.000}$ $0.341_{\pm 0.072}$ $0.263_{\pm 0.081}$
CALB MAPE $0.140_{\pm 0.009}$ $1.536_{\pm 0.147}$ $0.146_{\pm 0.051}$ $0.256_{\pm 0.070}$
15%-ACC $0.704_{\pm 0.053}$ $0.219_{\pm 0.109}$ $0.767_{\pm 0.088}$ $0.404_{\pm 0.052}$
@@ 4.6. Cross-domain Transferability (RQ4)
In real-world situations, batteries of new types are often fewer than batteries of old types so that it is important to develop a model that has cross-domain transferability against the domain shifts caused by battery types. In this work, we evaluate the cross-domain transferability of models by investigating if a model pretrained on Li-ion dataset can be transferred to other domains, and report the results in Table 6. Specifically, we study the transferability of CPMLP which is the best model in Li-ion domain. Three transferability evaluations are considered: (1) Frozen: Directly applying the pretrained CPMLP on target domains. (2) Fine-tuning (FT): Fine-tune all parameters of the pretrained CPMLP. (3) Domain adaptation (DA): we implement a domain adaptation technique by minimizing maximum mean discrepancy (Han et al., 2022) between the target-domain and source-domain embeddings produced by the last hidden layer. By expectation, with transfer learning techniques, an effective transferrable model should outperform the best baseline that is trained using only target-domain training data.
Table 6 shows that frozen pretrained models perform poorly on target domains due to significant domain shifts. While fine-tuning and domain adaptation improve performance over frozen models, they underperform the baseline in most cases. This suggests that extracting relevant information from diverse lab-test data (with hundreds of aging conditions) to other battery types remains challenging. The cross-domain evaluation results indicate that BatteryLife still has substantial room for improvements in cross-domain transferability.
@@ 5. CONCLUSION AND FUTURE WORK
6. ACKNOWLEDGMENTS
The authors acknowledge the financial support of the National Key R&D Program of China (No. 2023YFB2503600). This work is also supported by research grants from the National Natural Science Foundation of China (No. 52207230, No. 92372109 and No. 62206067), the Guangzhou-HKUST(GZ) Joint Funding Program (No. 2023A03J0003, No. 2023A03J0103 and No. 2023A03J0673), and the Guangzhou Municipal Science and Technology Project (No. 2024A04J4216).
@@ A.1. Data Preprocessing
For all datasets, the present study filtered out the batteries whose SOH is not degraded to $\lambda+2.5\%$ ($\lambda=90\%$ for the CALB dataset and $\lambda=80\%$ for other datasets), and then conducted the data statistics. Note that (1) BatteryArchive is an active website, and the data statistics were conducted on the available data until September 2024 in this paper. (2) Stanford (Cui et al., 2024) were released multiple times and the data used in this work is from the first release. As for BatteryLife, we further excluded the batteries whose life labels are no larger than 100. Additionally, we found that some batteries in the UL_PUR dataset contain sudden significant SOH drops without recovery. We conjecture that these drops are caused by equipment faults and thus removed these batteries. The data statistics were conducted on the remaining batteries. The statistics of BatteryLife were conducted on the remaining batteries.
Furthermore, some outliers might exist in the cycling data. For the 7 datasets standardized by BatteryML (Zhang et al., 2023a), we detected abnormal cycles by running a median filtering method on discharge capacity degradation trajectory following BatteryML (Zhang et al., 2023a) and removed the detected outliers. As for other datasets, we first removed the rest performance tests (RPT) and formation cycles. The RPT cycles are removed because they comply with charge/discharge protocols that are different from the cycling protocols, and the RPT cycles are also generally detected as outliers by BatteryML (Zhang et al., 2023a). Likewise, formation cycles are removed. Moreover, some cycles of sudden rise or drop in discharge capacities are manually removed. After data cleaning, each battery was stored in a data format consistent with BatteryML (Zhang et al., 2023a).
@@ A.3. Overview of The Integrated Datasets
  • HNEI (Devie et al., 2018): In this dataset, there are 14 18650-format lithium-ion batteries (LG Chem, ICR18650 C2) whose cathode is a mixture of $\rm LiCoO_{2}$ and $\rm LiNi_{0.4}Co_{0.4}Mn_{0.2}O_{2}$, and the anode is graphite. The electrolyte composition is not disclosed, and the nominal capacity is 2.8 Ah for all batteries. There are 2 different protocols with a working temperature of 25 degree Celsius.
@@
  • SNL (Preger et al., 2020): This dataset consists of 50 batteries in the format of 18650 cylindrical batteries. The positive electrode covers $\rm LiFePO_{4}$, $\rm LiNi_{0.81}Co_{0.14}Al_{0.05}O_{2}$, and $\rm LiNi_{0.84}Mn_{0.06}Co_{0.1}O_{2}$, while the negative electrode is graphite. The electrolyte formula is unknown. There are 3 nominal capacities consisting of 1.1 Ah, 3.2 Ah, and 3.0 Ah. There are 23 different protocols in this dataset with 3 different working temperatures (15°C, 25°C and 35°C).
@@
  • MICH_EXP (Mohtat et al., 2021): This dataset consists of 18 lithium-ion batteries in the format of pouch batteries. The positive electrode is $\rm LiNi_{1/3}Co_{1/3}Mn_{1/3}O_{2}$ (NCM111), and the negative electrode is $\rm graphite$. The electrolyte consists of 1M $\rm LiPF_{6}$ salt in solvents of ethylene carbonate (EC) and ethyl methyl carbonate (EMC) with a ratio of 3:7. The nominal capacity is 5.0 Ah. There are 6 different protocols with 3 different working temperatures in total (-5°C, 25°C, and 45°C).
@@
  • Tongji (Zhu et al., 2022): This dataset contains 108 commercial Li-ion batteries in a format of 18650 cylindrical batteries, covering 3 chemical systems consisting of 3 different positive electrodes, 2 different negative electrodes, and an identical electrolyte. This dataset includes 2 nominal capacities (3.5 Ah and 2.5 Ah), 3 operating temperatures, and 6 charge/discharge protocols.
  • Stanford (Cui et al., 2024): This dataset covers 41 batteries in the format of pouch batteries whose positive electrode is $\rm LiNi_{0.5}Mn_{0.3}Co_{0.2}O_{2}$, negative electrode is artificial graphite. The electrolyte is a mixture of $\rm LiPF_{6}$, ethylene carbonate, ethyl methyl carbonate, dimethyl carbonate, and vinylene carbonate. The nominal capacity of these batteries is 0.24 Ah. All batteries underwent different formation protocols but were cycled with an identical charge/discharge protocol under a working temperature of 30 degree Celsius.
  • XJTU (Wang et al., 2024c): There are 23 batteries in this dataset in the format of 18650 cylindrical batteries. For all batteries, the positive electrode is $\rm LiNi_{0.5}Co_{0.2}Mn_{0.3}O_{2}$ and the negative electrode is graphite. The electrolyte is unknown. The nominal capacity is 2 Ah. This dataset includes 2 different operation conditions with a working temperature of 20 degree Celsius.
  • ISU_ILCC (Li et al., 2024): This dataset comprises 240 502030-size Li-polymer batteries. The positive electrode is lithium nickel manganese cobalt oxide, and the negative electrode is graphite. The electrolyte remains unknown. And the nominal capacity is 0.25 Ah. There are 202 different charge/discharge protocols in this dataset, with 30 degree Celsius as the working temperature.
  • Zn-ion: This dataset consists of 100 zinc-ion batteries in the format of coin batteries. The positive electrode is $\rm MnO_{2}$ while the negative electrode is zinc metal. The electrolyte is unknown. The discharge capacity of the $10^{th}$ cycle was taken as the nominal capacity for each battery. All batteries underwent the same charge/discharge protocol at 3 different working temperatures.
@@ B.1. Further Details on Hyperparameters
For each model, we tried more than 20 different sets of hyperparameters and selected the hyperparameters that led to the lowest MAPE on the validation sets. The search ranges of hyperparameters are described as follows: batch size was chosen from [16, 32, 64, 128], learning rate was selected from [$5\times 10^{-4}$, $5\times 10^{-3}$, $1\times 10^{-3}$, dropout rate was chosen from [0, 0.05, 0.1], and embedding dimension was selected from [32, 64, 128, 256]. Additionally, for CyclePatch models, both the number of intra-cycle encoder layers and the number of inter-cycle encoder layers were tuned from 0 to 12. The number of CNN layers is tuned from 1 to 8.
@@ B.2. Implementation Details of Baselines
In this subsection, we introduce how we implemented the baselines in detail. First, the cycling data are resampled into $\textbf{X}\in\mathbb{R}^{3\times 30000}$ according to the process introduced in Section 4.1. After that, the three variables in each cycle are normalized as follows:
@@
where $\textbf{p}_{i}\in\mathbb{R}^{300}$, $\textbf{v}_{i}\in\mathbb{R}^{300}$, $\textbf{s}_{i}\in\mathbb{R}^{300}$ represent the capacity, voltage, and current records in the $i^{th}$ cycle. The $Q_{nominal}$ denotes the nominal capacity. Let $\overline{\textbf{X}}_{i}$ represent the normalized cycling data of the $i^{th}$ cycle. Then, Transformer encoder, LSTM, BiLSTM, GRU, BiGRU can directly model $\overline{\textbf{X}}_{i}$ and produce prediction as follows:
@@
As for the DLinear, PatchTST, Autoformer, iTransformer, and MICN, we use the implementation in https://github.com/thuml/Time-Series-Library. Specifically, we slightly revise the classification implementations to make models usable for BLP.